# KernelSHAP-IQ: Weighted Least Square Optimization for Shapley Interactions

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Potential Question: “Does the location of my property affect its price $\hat{y}=4.54$ (in 100k\$)?
@@
Both, SII and $1$-SII are the SV for individuals. In this case, besides the classical representation, it is well-known that the SV yields an optimal $1$-additive approximation of the cooperative game, which minimizes the weighted least square (WLS) loss over all subsets constrained on the efficiency property (Charnes et al., 1988). Extensions of such representations for interactions exist for constant weights, which yields the Banzhaf interaction index (Hammer & Holzman, 1992; Grabisch et al., 2000). For Shapley interactions, the Faithful Shapley Interaction Index (FSI) (Tsai et al., 2023) extends on the axiomatic structure of the least square family (Ruiz et al., 1998) by imposing Shapley-like axioms. However, while FSI yields a unique index, it neither yields SII nor $k$-SII and the dualism between both representations is not as clear as for the SV. In light of the exponential complexity of the SV and SII, which necessitates efficient approximation techniques, a WLS representation of SII is particularly desirable. In fact, KernelSHAP (Lundberg & Lee, 2017) directly utilizes this representation of the SV and yields state-of-the-art (SOTA) approximations in ML applications (Chen et al., 2023; Kolpaczki et al., 2023). For FSI, where a direct extension of KernelSHAP is possible, it has been shown that it is similarly superior over sampling-based approaches (Fumagalli et al., 2023; Kolpaczki et al., 2024).
Contribution. In this work, we show that SII is the solution to a WLS problem, which yields the optimal $k$-additive approximation of any cooperative game via $k$-SII and SII. We then present two variants of KernelSHAP Interaction Quantification (KernelSHAP-IQ), a direct extension of KernelSHAP for higher-order SII. We empirically demonstrate that KernelSHAP-IQ outperforms existing baselines and yields SOTA performance across various datasets and model classes. Our main contributions include:
@@
Related Work. In cooperative game theory, the studied set functions are known as transferable utility (TU) games, and $k$-additive TU-games, if $k$ is the maximum order (Grabisch, 2016). Extensions of the SV to interactions were proposed with SII (Grabisch & Roubens, 1999), FSI (Tsai et al., 2023) among others (Marichal & Roubens, 1999; Sundararajan et al., 2020). Recent work was largely motivated by the lack of efficiency of SII, which was recently resolved by the SII-based aggregation $k$-SII (Bordt & von Luxburg, 2023) that extends on (pairwise) Shapley interactions (Lundberg et al., 2020). The dualism between semivalues, such as the SV or Banzhaf value (Banzhaf III, 1964), and optimal WLS approximations is well known (Ruiz et al., 1998). For interactions, these results extend to Banzhaf interactions (Grabisch & Roubens, 1999; Grabisch et al., 2000). FSI provides a novel index, which directly yields this dualism for Shapley interactions. However, FSI differs from SII and $k$-SII, and the classical representation, as well as the link between orders remains mostly unknown (Tsai et al., 2023). Lastly, our approach can be viewed as a decomposition of the Shapley residuals (Kumar et al., 2021).
Approximations of Shapley interactions, were introduced for particular interaction indices by Sundararajan et al. (2020); Tsai et al. (2023) and arbitrary cardinal interaction indices (Fujimoto et al., 2006) by Fumagalli et al. (2023); Kolpaczki et al. (2024). They directly extend on methods for the SV (Castro et al., 2009; Covert & Lee, 2021; Kolpaczki et al., 2023). KernelSHAP (Lundberg & Lee, 2017) was introduced for the SV and extended to FSI (Tsai et al., 2023). Including SII in KernelSHAP also improved approximations of the SV (Pelegrina et al., 2023). Above approximations do not make any assumption about the game and were used for model-agnostic interpretability. For tree-based models, it was shown that SII can be computed much more efficiently (Muschalik et al., 2024; Zern et al., 2023), which extends TreeSHAP (Lundberg et al., 2020; Yu et al., 2022) to SII.
@@ 2 From Shapley Values to Interactions
We consider a set of players $N:=\{1,\dots,n\}$ and the payout of a cooperative game $\nu:\mathcal{P}(N)\to\mathbb{R}$, where $\nu$ models the payout given a subset of players $T\subseteq N$. In the context of local interpretability, $N$ might be chosen as the set of features and $\nu$ the prediction of the model given only a subset of features. For simplicity, we consider mean-centered games requiring $\nu(\emptyset)=0$. In the following, we are interested in quantifying the contribution of a set of players $S\subseteq N$ to the payout $\nu$. For local interpretability, this is known as feature attribution for individual players and feature interaction for multiple players. For individual players, the fair attribution is the well-known Shapley value (SV) (Shapley, 1953)
$\mathcal{\phi}^{\text{SV}}(i)=\sum\nolimits_{T\subseteq N\setminus i}\tfrac{(n-1-t)!\cdot t!}{n!}\underbrace{\left[\nu(T\cup i)-\nu(T)\right]}_{=:\Delta_{i}(T)}.$
The SV is the unique attribution measure that fulfills the linearity, symmetry, dummy and in particular the efficiency axiom, which distributes the payout $\nu(N)-\nu(\emptyset)$ among the players, e.g. the prediction of a ML model. The SV is the weighted average over all marginal contributions $\Delta_{i}(T):=\nu(T\cup i)-\nu(T)$. For two players $ij$ the marginal contribution to $T\subseteq N\setminus ij$ can be extended using the recursion
@@
i.e. the effect of adding both players $ij$, minus their individual contributions. For feature interactions, a positive value indicates meaningful joint information, e.g. exact position versus latitude and longitude separately, whereas a negative value indicates redundancy. Using the above intuition, the marginal contributions can be extended to any subset, which is known as the discrete derivative $\Delta_{S}(T):=\sum_{L\subseteq S}(-1)^{s-l}\nu(T\cup L)$ of a set of players $S\subseteq N$ to a set of participating players $T\subseteq N\setminus S$. Similar to the SV, an interaction index is then defined as an weighted average over discrete derivatives.
@@
$\phi^{\text{SII}}(S):=\sum\nolimits_{T\subseteq N\setminus S}\frac{(n-s-t)!\cdot t!}{(n-s+1)!}\Delta_{S}(T).$
@@
The SII includes the SV as $\phi^{\text{SII}}_{1}=\phi^{\text{SV}}$. It was shown by Grabisch & Roubens (1999) that the SII satisfies the (generalized) linearity, symmetry and dummy axiom. The SII is further the unique interaction index that additionally satisfies the recursive axiom, which naturally extends the recursion of discrete derivatives and links higher- to lower-order interactions. For pairwise interaction, it requires
@@ 2.1 $k$-Additive Interactions and $k$-Shapley Values
Note that 1-SII is again the SV, i.e. $\Phi_{1}\equiv\phi^{\text{SV}}$. Furthermore, $k$-SII requires SII values up to order $k$ and the highest order is equal to SII. It was shown that $\Phi_{k}$ is efficient and that the Bernoulli numbers are the unique choice of weighting to achieve this property (Bordt & von Luxburg, 2023). For $k=n$ the explicit form is drastically simplified and corresponds to the well known Möbius transform (Bordt & von Luxburg, 2023). The $k$-SII therefore defines a flexible low-complexity $k$-additive interaction index, which is a straight-forward aggregation of SII and can be directly used for an interpretable representation. In Section 3, we will show and leverage that $k$-SII is directly linked to an optimal $k$-additive approximation of $\nu$.
@@ 2.2 Approximations of Shapley Interactions
In applications, the limiting factor of SII and the SV are the amount of evaluations of $\nu$, where both require $2^{n}$ evaluations for exact computation. Therefore, multiple approximation techniques have been proposed, as illustrated in Table 1. In the following, we introduce the main approaches for SII, which are based on Monte Carlo sampling of different representations, and extend on techniques for the SV. These methods will be used as baseline methods in our experiments. Permutation Sampling (Castro et al., 2009; Tsai et al., 2023) samples random permutations $\pi$ of $N$ and is based on the representation $\phi^{\text{SII}}(S)=\mathbb{E}_{\pi}[\mathbf{1}_{\pi\in S}\Delta_{S}(u_{S}(\pi))]$, where $u_{S}(\pi)$ is the set of elements that appear in $\pi$ before any element of $S$, and $\mathbf{1}_{\pi\in S}$ is one, if all elements of $S$ appear consecutively in $\pi$, and zero otherwise. For permutation sampling only a small fraction of interaction estimates are updated with each evaluation. An alternative method is SHAP-IQ (Fumagalli et al., 2023) that samples subsets $T\subseteq N$ directly and utilizes a representation $\phi^{\text{SII}}(S)=\mathbb{E}_{T}[\nu(T)\omega_{s}(t,|T\cap S|)]$, where $\omega_{s}$ are SII-specific weights adjusted to the sampling distribution. SHAP-IQ can therefore utilize every evaluation to update all interaction estimates. SVARM-IQ (Kolpaczki et al., 2024) further improves approximation by stratifying over subset sizes $t$ and intersections $T\cap S=L$ with
@@
where $\tilde{\omega}_{s}(t,L):=\omega_{s}(t,\ell)\cdot p(|T|=t,T\cap S=L)$, i.e. adjusted by the strata probabilities. The core idea is to compute strata estimates by sampling $T\subseteq N$ and treat it as a random sample for one specific stratum of every interaction. This allows again to utilize every evaluation for all interaction estimates. So far, it was shown empirically that SVARM-IQ is the most powerful estimator for SII (Kolpaczki et al., 2024).
@@ 2.3 Approximating Shapley Values with KernelSHAP
$\phi^{\text{SV}}$ $=\operatorname*{arg\,min}\nolimits_{\phi\in\mathbb{R}^{n}}\mathbb{E}_{T\sim p}\left[\left(\nu(T)-\sum\nolimits_{i\in T}\phi(i)\right)^{2}\right],$
such that $\text{such that }\sum\nolimits_{i\in N}\phi(i)=\nu(N),$
where $p(T)\propto\mu(t):=\binom{n-1}{t-2}^{-1}$ for $11\leq t\leq n-$ 1.
@@
Method SV SII ($k$-SII) Permutation (Castro et al., 2009) (Tsai et al., 2023) SHAP-IQ (Covert & Lee, 2021) (Fumagalli et al., 2023) SVARM (Kolpaczki et al., 2023) (Kolpaczki et al., 2024) KernelSHAP (Lundberg & Lee, 2017) (ours)
KernelSHAP first approximates the expectation using Monte Carlo sampling. Then, the optimization problem is solved exactly with the approximated objective. The constraint is thereby included with sets $T\in\{\emptyset,N\}$ and a large weight $\mu_{\infty}\gg 1$, which we rigorously prove later in Theorem 3.6. While the theoretical analysis of KernelSHAP remains difficult, it was shown that the approach works especially well for ML applications (Chen et al., 2023; Kolpaczki et al., 2023; Lundberg & Lee, 2017). In the following, we rigorously describe this approach and discover such a representation for SII, which we use for efficient approximation.
@@ 3 Weighted Least Square Optimization for Shapley Interactions
KernelSHAP utilizes that the SV is the solution to a WLS problem yielding an optimal $1$-additive approximation of $\nu$ with $\nu(T)\approx\sum_{i\in T}\phi^{\text{SV}}(i)$, which is exact for $\nu(N)=\sum_{i\in N}\phi^{\text{SV}}(i)$ due to the efficiency axiom. However, linking higher orders of SII to such an optimal approximation of $\nu$ has remained unknown. In order to extend KernelSHAP to interactions, recent research introduced the FSI as an alternative index (Tsai et al., 2023). FSI constructs an optimal $k$-additive approximation of $\nu$ by extending on the KernelSHAP representation with interactions up to order $k$. One disadvantage of FSI is that the solution directly depends on $k$, which complicates the relationship between representations of different order. Furthermore, in contrast to the SV, the exact representation of lower-order FSI with discrete derivatives remains unknown (Tsai et al., 2023). In contrast, we show that $k$-SII constructed from SII values yield an appealing iterative approximation of $\nu$, which additively extends on the previous approximation by including a weighted average of SII of order $k$. We then show that SII via $k$-SII is linked to an optimal $k$-additive approximation of $\nu$. This yields a suitable framework, where SII is fixed independent of $k$ and $k$-SII yields an interpretable $k$-additive interaction index, as summarized in Figure 3. Utilizing these results, we introduce KernelSHAP-IQ, an extension of KernelSHAP for Shapley interactions. We rigorously prove its convergence for pairwise interactions and conclude with empirically validated conjectures for higher orders. All proofs are deferred to Appendix A.

### v2

@@ Abstract
Shapley Interactions, Shapley Value, Explainable Artificial Intelligence, Game Theory
@@ 1 Introduction
Potential Question: “Does the location of my property affect its price $\hat{y}=4.54$ (in 100k\$)?”
@@
Both, SII and $1$-SII are the SV for individuals. In this case, besides the classical representation, it is well-known that the SV yields an optimal $1$-additive approximation of the cooperative game, which minimizes the weighted least square (WLS) loss over all subsets constrained on the efficiency property (Charnes et al., 1988). Extensions of such representations for interactions exist for constant weights, which yields the Banzhaf interaction index (Hammer & Holzman, 1992; Grabisch et al., 2000). For Shapley interactions, the Faithful Shapley Interaction Index (FSI) (Tsai et al., 2023) extends on the axiomatic structure of the least square family (Ruiz et al., 1998) by imposing Shapley-like axioms. However, while FSI yields a unique index, it neither yields SII nor $k$-SII and the dualism between both representations is not as clear as for the SV. In light of the exponential complexity of the SV and SII, which necessitates efficient approximation techniques, a WLS representation of SII is particularly desirable. In fact, KernelSHAP (Lundberg & Lee, 2017) directly utilizes this representation of the SV and yields state-of-the-art (SOTA) approximations in ML applications (Chen et al., 2023; Kolpaczki et al., 2024a). For FSI, where a direct extension of KernelSHAP is possible, it has been shown that it is similarly superior over sampling-based approaches (Fumagalli et al., 2023; Kolpaczki et al., 2024b).
Contribution. In this work, we show that SII is the solution to a WLS problem, which yields the optimal $k$-additive approximation of any cooperative game via $k$-SII and SII. We then present two variants of KernelSHAP Interaction Quantification (KernelSHAP-IQ), a direct extension of KernelSHAP for higher-order SII. We empirically demonstrate that KernelSHAP-IQ outperforms existing baselines and yields SOTA performance across various datasets and model classes. Therein, our main contributions include:
@@
Related Work. In cooperative game theory, the studied set functions are known as transferable utility (TU) games, and $k$-additive TU-games, if $k$ is the maximum order (Grabisch, 2016). Extensions of the SV to interactions were proposed with SII (Grabisch & Roubens, 1999), FSI (Tsai et al., 2023) among others (Marichal & Roubens, 1999; Sundararajan et al., 2020). Recent work was largely motivated by the lack of efficiency of SII, which was recently resolved by the SII-based aggregation $k$-SII (Bordt & von Luxburg, 2023) that extends on (pairwise) Shapley interactions (Lundberg et al., 2020). The dualism between semivalues, such as the SV or Banzhaf value (Banzhaf III, 1964), and optimal WLS approximations is well known (Ruiz et al., 1998). For interactions, these results extend to Banzhaf interactions (Grabisch & Roubens, 1999; Grabisch et al., 2000). FSI provides a novel index, which directly yields this dualism for Shapley interactions. However, FSI differs from SII and $k$-SII, and the classical representation, as well as the link between orders remains mostly unknown (Tsai et al., 2023). Lastly, our approach is linked to the Shapley residuals (Kumar et al., 2021).
Approximations of Shapley interactions, were introduced for particular interaction indices by Sundararajan et al. (2020); Tsai et al. (2023) and arbitrary cardinal interaction indices (Fujimoto et al., 2006) by Fumagalli et al. (2023); Kolpaczki et al. (2024b). They directly extend on methods for the SV (Castro et al., 2009; Covert & Lee, 2021; Kolpaczki et al., 2024a). KernelSHAP (Lundberg & Lee, 2017) was introduced for the SV and extended to FSI (Tsai et al., 2023). Including SII in KernelSHAP also improved approximations of the SV (Pelegrina et al., 2023). Above approximations do not make any assumption about the game and were used for model-agnostic interpretability. For tree-based models, it was shown that SII can be computed much more efficiently (Muschalik et al., 2024; Zern et al., 2023), which extends TreeSHAP (Lundberg et al., 2020; Yu et al., 2022) to SII.
@@ 2 From Shapley Values to Interactions
We consider a set of players $N:=\{1,\dots,n\}$ and the payout of a cooperative game $\nu:\mathcal{P}(N)\to\mathbb{R}$, where $\nu$ models the payout given a subset of players $T\subseteq N$. In the context of local interpretability, $N$ might be chosen as the set of features and $\nu$ the prediction of the model given only a subset of features. For simplicity, we consider games with $\nu(\emptyset)=0$. In the following, we are interested in quantifying the contribution of a set of players $S\subseteq N$ to the payout $\nu$. For local interpretability, this is known as feature attribution for individual players and feature interaction for multiple players. For individual players, the fair attribution is the well-known Shapley value (SV) (Shapley, 1953)
$\mathcal{\phi}^{\text{SV}}(i)=\sum_{T\subseteq N\setminus i}\tfrac{(n-1-t)!\cdot t!}{n!}\underbrace{\left[\nu(T\cup i)-\nu(T)\right]}_{=:\Delta_{i}(T)}.$
The SV is the unique attribution measure that fulfills the linearity, symmetry, dummy, and in particular the efficiency axiom, which distributes the payout $\nu(N)-\nu(\emptyset)$ among the players, e.g. the prediction of a ML model. The SV of player $i$ is the weighted average over all marginal contributions $\Delta_{i}(T):=\nu(T~{}\cup~{}i)-\nu(T)$. For two players $ij$ the marginal contribution to $T\subseteq N\setminus ij$ can be extended using the recursion
@@
i.e. the effect of adding both players $ij$, minus their individual contributions. For feature interactions, a positive value indicates meaningful joint information, e.g. exact position versus latitude and longitude separately, whereas a negative value indicates redundancy. Using above intuition, the marginal contributions can be extended to any subset, which is known as the discrete derivative $\Delta_{S}(T):=\sum_{L\subseteq S}(-1)^{s-l}\nu(T\cup L)$ of a set of players $S\subseteq N$ to a set of participating players $T\subseteq N\setminus S$. Similar to the SV, an interaction index is then defined as a weighted average over discrete derivatives.
@@
$\phi^{\text{SII}}(S):=\sum_{T\subseteq N\setminus S}\frac{(n-s-t)!\cdot t!}{(n-s+1)!}\Delta_{S}(T).$
@@
The SII includes the SV as $\phi^{\text{SII}}_{1}\equiv\phi^{\text{SV}}$. It was shown by Grabisch & Roubens (1999) that the SII satisfies the (generalized) linearity, symmetry and dummy axiom. The SII is further the unique interaction index that additionally satisfies the recursive axiom, which naturally extends the recursion of discrete derivatives and links higher- to lower-order interactions. For pairwise interaction, it requires
@@ 2.1 $k$-Additive Interactions and $k$-Shapley Values
Note that 1-SII is again the SV, i.e. $\Phi_{1}\equiv\phi^{\text{SV}}$. Furthermore, $k$-SII requires SII values up to order $k$ and the highest order is equal to SII. It was shown that $\Phi_{k}$ is efficient and that the Bernoulli numbers are the unique choice of weighting to achieve this property (Bordt & von Luxburg, 2023). For $k=n$ the explicit form is drastically simplified and corresponds to the well-known Möbius transform (Bordt & von Luxburg, 2023). The $k$-SII therefore defines a flexible low-complexity $k$-additive interaction index, which is a straight-forward aggregation of SII and can be directly used for an interpretable representation. In Section 3, we will show and leverage that $k$-SII is directly linked to an optimal $k$-additive approximation of $\nu$.
@@ 2.2 Approximations of Shapley Interactions
In applications, the limiting factor of SII and the SV are the amount of evaluations of $\nu$, where both require $2^{n}$ evaluations for exact computation. Therefore, multiple approximation techniques have been proposed, as illustrated in Table 1. In the following, we introduce the main approaches for SII, which are based on Monte Carlo sampling of different representations, and extend on techniques for the SV. These methods will be used as baseline methods in our experiments. Permutation Sampling (Castro et al., 2009; Tsai et al., 2023) samples random permutations $\pi$ of $N$ and is based on the representation $\phi^{\text{SII}}(S)=\mathbb{E}_{\pi}[\mathbf{1}_{\pi\in S}\Delta_{S}(u_{S}(\pi))]$, where $u_{S}(\pi)$ is the set of elements that appear in $\pi$ before any element of $S$, and $\mathbf{1}_{\pi\in S}$ is one, if all elements of $S$ appear consecutively in $\pi$, and zero otherwise. For permutation sampling only a small fraction of interaction estimates are updated with each evaluation. An alternative method is SHAP-IQ (Fumagalli et al., 2023) that samples subsets $T\subseteq N$ directly and utilizes a representation $\phi^{\text{SII}}(S)=\mathbb{E}_{T}[\nu(T)\omega_{s}(t,|T\cap S|)]$, where $\omega_{s}$ are SII-specific weights adjusted to the sampling distribution. SHAP-IQ can therefore utilize every evaluation to update all interaction estimates. SVARM-IQ (Kolpaczki et al., 2024b) further improves approximation by stratifying over subset sizes $t$ and intersections $T\cap S=L$ with
@@
where $\tilde{\omega}_{s}(t,L):=\omega_{s}(t,\ell)\cdot p(|T|=t,T\cap S=L)$, i.e. adjusted by the strata probabilities. The core idea is to compute strata estimates by sampling $T\subseteq N$ and treat it as a random sample for one specific stratum of every interaction. This allows again to utilize every evaluation for all interaction estimates. It was shown empirically that SVARM-IQ is the most powerful estimator for SII (Kolpaczki et al., 2024b).
@@ 2.3 Approximating Shapley Values with KernelSHAP
$\phi^{\text{SV}}&=\operatorname*{arg\,min}_{\phi\in\mathbb{R}^{n}}\sum_{T\subseteq N}^{0<|T|<n}\mu(t)\left[\nu(T)-\sum_{i\in T}\phi(i)\right]^{2},\\ &\text{such that }\sum_{i\in N}\phi(i)=\nu(N),\end{aligned}$ (1)
where $p(T)\propto\mu(t):=\binom{n-2}{t-1}^{-1}$ for $11\leq t\leq n-$ 1.
@@
Method SV SII ($k$-SII) Permutation (Castro et al., 2009) (Tsai et al., 2023) SHAP-IQ (Covert & Lee, 2021) (Fumagalli et al., 2023) SVARM (Kolpaczki et al., 2024a) (Kolpaczki et al., 2024b) KernelSHAP (Lundberg & Lee, 2017) KernelSHAP-IQ
KernelSHAP considers above objective as an expectation
$\sum_{T\subseteq N}^{0<t<n}\mu(t)[\nu(T)-\sum_{i\in T}\phi(i)]^{2}=\mathbb{E}_{T}[(\nu(T)-\sum_{i\in T}\phi(i))^{2}]$
over sampled subsets $T\sim p(T)\propto\mu(t)$, and approximates the expectation using Monte Carlo sampling. Then, the WLS problem is solved with the approximated objective. The constraint is thereby included with sets $T\in\{\emptyset,N\}$ and a large weight $\mu_{\infty}\gg 1$, where we prove in Theorem 3.6 that this yields the SV. While the theoretical analysis of KernelSHAP remains difficult, it was shown that the approach works especially well for ML applications (Chen et al., 2023; Kolpaczki et al., 2024a; Lundberg & Lee, 2017). In the following, we rigorously describe this approach and discover such a representation for SII, which we use for efficient approximation.
@@ 3 Weighted Least Square Optimization for Shapley Interactions
The SII is an axiomatic extension of the SV and constructs $k$-SII as an efficient $k$-additive interaction index used for interpretability. In this section, we give an alternative characterization of SII, as the solution to a WLS problem, which extends on the representation of the SV given in Equation 1. Based on this representation, we introduce KernelSHAP-IQ, an extension of KernelSHAP for Shapley interactions. We rigorously prove its convergence for pairwise interactions and conclude with empirically validated conjectures for higher orders. All proofs are deferred to Appendix A.
The representation of the SV in Equation 1, utilized by KernelSHAP, shows that the approximation $\sum_{i\in T}\phi(i)\approx\nu(T)$ of the game is optimal in terms of the WLS objective, if and only if $\phi(i)$ are the SVs. This approximation can be viewed as a $1$-additive approximation of the game, which we now directly extend to $k$-additive approximations.
Definition 3.1 ($k$-additive Approximation).
The $k$-additive approximation induced by the $k$-SII $\Phi_{k}$ is given by
$\hat{\nu}_{k}(T):=\sum_{S\subseteq T}^{1\leq|S|\leq k}\Phi_{k}(S)\text{ for }T\subseteq N.$
Clearly, $\hat{\nu}_{1}$ is the approximation given in Equation 1. Linking higher orders of the SII to such an optimal approximation of $\nu$ has remained unknown. In order to extend KernelSHAP to interactions, recent research introduced the FSI as an alternative index (Tsai et al., 2023). FSI constructs an optimal $k$-additive approximation of $\nu$ by extending on the KernelSHAP representation with interactions up to order $k$. One disadvantage of FSI is that the solution directly depends on $k$, which complicates the relationship between representations of different order. Furthermore, the exact representation of lower-order FSI with discrete derivatives remains unknown (Tsai et al., 2023). In contrast, we show that $k$-SII constructed from SII values yield an appealing iterative approximation of $\nu$, which additively extends on the previous approximation by including a weighted average of SII of order $k$. We then show that SII via $k$-SII is linked to an optimal $k$-additive approximation of $\nu$. This yields a suitable framework, where SII is fixed independent of $k$ and $k$-SII yields an interpretable $k$-additive interaction index, as summarized in Figure 3.
3.1 Optimal Approximations via Shapley Interactions